# Mean-field games with rough common noise I: the linear-quadratic case

### v1

@@ Abstract
Motivated by mean-field games (MFG) with common noise on the one hand and pathwise stochastic control theory on the other, we formulate here a linear-quadratic (LQ) MFG with rough common noise, along with a satisfactory well-posedness theory for the linear-quadratic case. A novel Volterra-type (or mild) formulation allows to keep technical (rough-stochastic) consideration to a minimum. We derive a characterization of the optimal state and optimal control through a rough forward-backward SDE (rough FBSDE), and provide an existence and uniqueness result under the usual assumptions. Our theory is accompanied by stability estimates with respect to initial data and common noise while we also establish continuity of what we call the Itô-Lions-Lyons map for rough mean-field games. Finally, we discuss randomization of the rough common noise under appropriate conditions on the coefficients. When the latter is given by the Stratonovich lift of a Brownian motion independent of the idiosyncratic noise, we show that solutions of the rough LQ MFG coincide with those obtained by conditioning on the common noise.
@@ 1.1 Problem formulation
$\mathcal{J}(\alpha;\mu,\xi;\omega):=\mathbb{E}\Big[\int_{0}^{T}\ell(t,X^{\alpha,\mu,\xi}_{t},\mu_{t},\alpha_{t})\mathrm{d}t+\Phi(X^{\alpha,\mu,\xi}_{T},\mu_{T})\Big|\mathcal{F}^{\perp}_{T}\Big]\hskip 0.0pt,\\$
@@
$\mu_{t}=\mathcal{L}\left(X_{t}^{\alpha^{\ast}(\mu),\mu,\xi}\hskip 1.69998pt|\hskip 1.69998pt\mathcal{F}^{{\perp}}_{T}\right).$
@@
$\mathcal{L}\left(\overline{X}^{\overline{\alpha}}_{t}\hskip 1.69998pt|\hskip 1.69998pt\mathcal{F}^{{\perp}}_{T}\right)=\mathcal{L}\left({X}^{\boldsymbol{\eta},{\alpha}(\boldsymbol{\eta}),\mu(\boldsymbol{\eta}),\xi}_{t}\right)\Big|_{\boldsymbol{\eta}={W}^{\perp}(\omega^{\prime\prime})}.$
@@ 1.2 Main results
$+\int_{0}^{t}\big(A^{1}_{s}X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{s}+C^{1}_{s}m_{s}\big)\mathrm{d}\boldsymbol{\eta}_{s},$
where $(m,m^{\prime})$ is a fixed controlled rough path and $C^{1}m\mathrm{d}\boldsymbol{\eta}$ a meaningful (Lyons-Gubinelli) rough integral (at first reading, assume $A^{1},C^{1}$ may be assumed constant in time, but we also treat controlled time-dependence, with Gubinelli derivatives $(A^{1})^{\prime},(C^{1})^{\prime}$). The agent aims to minimize the quadratic cost functional
$\mathcal{J}(\alpha)=\mathbb{E}\!\bigg[\int_{0}^{T}\frac{1}{2}\Big((X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t})^{\top}Q_{t}X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}+\alpha_{t}^{\top}R_{t}\alpha_{t}\\ +(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}-S_{t}m_{t})^{\top}\bar{Q}_{t}(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}-S_{t}m_{t})\Big)\mathrm{d}t\\ +\frac{1}{2}\Big(\big(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{T}\big)^{\top}QX^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{T}+\big(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{T}-Sm_{T}\big)^{\top}\bar{Q}\big(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{T}-Sm_{T}\big)\Big)\bigg].$ (2)
@@
It is customary in LQ MFG theory, that one does not try to solve the fixed point problem with respect to the flow of probability measures directly, but rather solves the adjoint fixed point problem with respect to the corresponding mean flows. Of course the two fixed point problems are equivalent, if the initial distribution of the representative agent is Gaussian. In our setting, the mean flow formulation is more beneficial since a) it leads to a natural extension of the classical theory to a rough setting and b) it does not make use of measure-valued paths/flows and avoids associated technical considerations.
Our results can be seen as extensions of [19, Thm 3.34], also Proposition 3.35 therein, on global solutions. A novel feature or our analysis is to exploit the affine-linear rough noise structure of the agent dynamics (1) to rewrite it in a mild/Volterra form. Our first main result, stated here in a simplified form, establishes the equivalence between Volterra and rough MFGs; the precise formulation and proof follows from the combination of Theorems 4.5, 4.8 and 4.12 below.
@@
$\left\{\begin{aligned} &{X}^{m,\xi}_{t}=\mathcal{K}(t,0)\xi+M(m)_{t}\\ &\quad\quad+\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}{X}^{m,\xi}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}Y^{m,\xi}_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}(t,s)\Sigma_{s}\mathrm{d}W_{s},\\ &Y^{m,\xi}_{t}=\mathcal{K}(T,t)^{\top}\big((Q+\bar{Q})X_{T}^{m,\xi}-\bar{Q}Sm_{T}\big)\\ &\quad\quad+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{m,\xi}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{m,\xi}_{s}\Big)\mathrm{d}s-\int_{t}^{T}\mathcal{K}(s,t)^{\top}Z^{m,\xi}_{s}\mathrm{d}W_{s}.\end{aligned}\right$.
@@
A detailed version of the this theorem is contained in Section 5 below as Theorem 5.3.
Finally, Section 6 is devoted to randomization, already discussed in the end of Section $1.$
@@ 1.3 Structure of the paper
After discussing frequently-used notation and some preliminaries on rough paths and rough (stochastic) integration at Section 2, we begin our approach to rough LQ MFG at Section 3. There, we develop the stochastic maximum principle in a Volterra setting by avoiding completely any mention of rough paths. This approach allows us to avoid calculations that would otherwise require strong moment conditions that in general are not possible to verify (at least in the context of Hölder rough paths). Then, at Section 4 we connect the Volterra and rough settings by delicate use of Duhamel’s principle, rigorously proved with a rough product formula. Thus, we obtain the aforementioned existence and uniqueness results for the Volterra and Rough FBSDEs, while also reducing the problem of solutions for the Volterra and rough LQ MFG problems to their deterministic counterparts. Next, we solve the deterministic forward-backward rough differential equation (FBRDE) by following the classical approach and imposing more general assumptions to account for the presence of the rough path. In Section 5, we establish stability estimates and local-Lipschitz continuity of the Itô-Lions-Lyons map with respect to rough common noise and initial conditions. Finally, Section 6 is devoted to randomization of the common noise and connection to the (conditional) stochastic problem that was discussed at the end of Section 1.1 . Auxiliary results on linear rough SDEs, accompanied by short proofs, are relegated to Appendix A.
@@ 2.2 Rough path theory
    The space of $\gamma$-Hölder continuous rough paths is denoted $\mathscr{C}^{\gamma}([0,T];\mathbb{R}^{p})$ and is equipped with the following seminorm and rough path metric: for any $\boldsymbol{\eta}=(\eta,\eta^{(2)}),\bar{\boldsymbol{\eta}}=(\bar{\eta},{\bar{\eta}}^{(2)})\in\mathscr{C}^{\gamma}([0,T];\mathbb{R}^{p}),$
@@
We shall always endow $\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$with the rough path metric $\rho_{\gamma}$. Moreover, we shall frequently write $\rho_{\gamma}(\boldsymbol{\eta})$ instead of $\rho_{\gamma}(\boldsymbol{\eta},0)$.
@@ 3. $(\text{iii})$
    In this case we write $(Z,Z^{\prime})\in\textbf{D}_{\eta}^{\beta,\beta^{\prime}}\mathbb{L}^{k}([0,T];V)$ and $(Z,Z^{\prime})\in\mathbf{D}_{\eta}^{2\beta}\mathbb{L}^{k}([0,T];V)$ if $\beta=\beta^{\prime}$. We equip the space $\mathbf{D}_{\eta}^{\beta,\beta^{\prime}}L^{k}$ with the norm
    $\llbracket(Z,Z^{\prime})\rrbracket_{\eta,\beta,\beta^{\prime},k}:=\left\lVert Z_{0}\right\rVert_{k}+\lVert\delta Z\rVert_{\beta,k}+\left\lVert Z^{\prime}_{0}\right\rVert_{k}+\lVert\delta Z^{\prime}\rVert_{\beta^{\prime},k}+\lVert\mathbb{E}_{\bullet}R^{Z}\rVert_{\beta+\beta^{\prime},k}.$
@@
    $\llbracket Z,Z^{\prime};\bar{Z},\bar{Z}^{\prime}\rrbracket_{\eta,\bar{\eta},\beta,\beta^{\prime},k}:=\|\delta(Z-\bar{Z})\|_{\beta,k}+\|\delta(Z^{\prime}-\bar{Z}^{\prime})\|_{\beta^{\prime},k}+\|Z_{0}-\bar{Z}_{0}\|_{k}$
@@
    $\|Z,Z^{\prime};\bar{Z},\bar{Z}^{\prime}\|_{\eta,\bar{\eta},\beta,\beta^{\prime},k}$ $\penalty 10000\ :=\|\delta(Z-\bar{Z})\|_{\beta,k}+\|\delta(Z^{\prime}-\bar{Z}^{\prime})\|_{\beta^{\prime},k}$
@@
We emphasize that in the case where $(Z,Z^{\prime})\in\mathbf{D}_{{\eta}}^{\beta,\beta^{\prime}}\mathbb{L}^{k}$ is deterministic then immediately we get $(Z,Z^{\prime})\in\mathscr{D}_{\eta}^{\beta,\beta^{\prime}}\mathbb{L}^{\infty}.$ The space of deterministic controlled rough paths is denoted by $\mathscr{D}_{\eta}^{\beta,\beta^{\prime}}$, or $\mathscr{D}_{\eta}^{\beta,\beta^{\prime}}([0,T];V)$ when we need to emphasize the codomain $V.$ Furthermore, in the deterministic setting we abbreviate the notation and we simply write $\|\cdot,\cdot;\cdot,\cdot\|_{\eta,\bar{\eta},\beta,\beta^{\prime}}$.
@@
$\mathop{\mathrm{ess\,sup}}_{t,\omega}|Z_{t}|\vee\|(Z,Z^{\prime})\|_{\boldsymbol{\eta},\beta,\beta^{\prime},\infty}\leq K\quad\mbox{and}\quad\|z_{0}\|_{k}\vee\|z^{\prime}_{0}\|_{k}\leq K_{0}$
@@
The next result is from [12, Thm 3.11] in case $\beta^{\prime}=\beta$, and [42, Thm 3.5] otherwise. It establishes well-posedness of (1).
@@
Let $p\in\mathbb{N},$ $\gamma\in(\frac{1}{3},\frac{1}{2})$ and $0<\beta^{\prime}\leq\beta\leq\gamma$ such that $\gamma+\beta+\beta^{\prime}>1$,
and $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$. Additionally, let $A,B,C:[0,T]\rightarrow\mathbb{R}^{d\times d}$ bounded and Borel measurable on $[0,T]$, $\big({A}^{1},(A^{1})^{\prime}\big),\big({C}^{1},(C^{1})^{\prime}\big)\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big),(m,m^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ deterministic controlled rough paths with respect to $\eta$}, $\xi\in\bigcap_{k=2}^{\infty}\mathbb{L}^{k}(\mathcal{F}_{0},\mathbb{R}^{d})$, and $\alpha(\omega,t),\Sigma(\omega,t)$ progressively measurable processes such that $\sup_{t\in[0,T]}\mathbb{E}[|\Sigma_{t}|^{k}]+\sup_{t\in[0,T]}\mathbb{E}[|\alpha_{t}|^{k}]<\infty$ for every $k\in[2,\infty)$. Then, there exists a unique solution $X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}$ of

### v2

# Mean-field games with rough common noise: the linear-quadratic case

@@ Abstract
Motivated by mean-field games (MFG) with common noise on the one hand and pathwise stochastic control theory on the other, we formulate here a linear-quadratic (LQ) MFG with rough common noise, along with a satisfactory well-posedness theory for the linear-quadratic case. A novel Volterra-type (or mild) formulation allows us to keep technical (rough-stochastic) considerations to a minimum. We derive a characterization of the optimal state and optimal control through a rough forward-backward SDE (rough FBSDE), and obtain existence and uniqueness. A number of stability estimates are established, and in particular we show continuous dependence of the Itô-Lions-Lyons map (equilibrium law as a function of the rough common noise). In a final section, we discuss independent Brownian randomization of the rough common noise. The resulting stochastic problem can be seen as an MFG extension of pathwise stochastic control problems.
@@ 1.1 Problem formulation
$\mathcal{J}(\alpha;\mu,\xi;\omega):=\mathbb{E}\Big[\int_{0}^{T}\ell(t,X^{\alpha,\mu,\xi}_{t},\mu_{t},\alpha_{t})\mathrm{d}t+\Phi(X^{\alpha,\mu,\xi}_{T},\mu_{T})\Big|\mathcal{F}^{\perp}_{T}\Big]\hskip 0.0pt,$
@@
$\mu_{t}=\mathcal{L}\big(X_{t}^{\alpha^{\ast}(\mu),\mu,\xi}\hskip 1.69998pt\big|\hskip 1.69998pt\mathcal{F}^{{\perp}}_{T}\big).$
@@
$\mathcal{L}\Big(\overline{X}^{\overline{\alpha}}_{t}\hskip 1.69998pt\Big|\hskip 1.69998pt\mathcal{F}^{{\perp}}_{T}\Big)=\mathcal{L}\left({X}^{\boldsymbol{\eta},{\alpha}(\boldsymbol{\eta}),\mu(\boldsymbol{\eta}),\xi}_{t}\right)\Big|_{\boldsymbol{\eta}={W}^{\perp}(\omega^{\prime\prime})}.$
@@ 1.2 Main results
$\quad+\int_{0}^{t}\big(A^{1}_{s}X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{s}+C^{1}_{s}m_{s}\big)\mathrm{d}\boldsymbol{\eta}_{s},$
where $(m,m^{\prime})$ is a fixed controlled rough path, later related to the mean of the process, cf. (3) below, and $C^{1}m\mathrm{d}\boldsymbol{\eta}$ a meaningful (classical) rough integral (at first reading, $A^{1},C^{1}$ may be assumed constant in time, but we also treat controlled time-dependence, with Gubinelli derivatives $(A^{1})^{\prime},(C^{1})^{\prime}$). The agent aims to minimize the quadratic cost functional
$\mathcal{J}(\alpha)=\mathbb{E}\!\bigg[\int_{0}^{T}\frac{1}{2}\Big((X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t})^{\top}Q_{t}X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}+\alpha_{t}^{\top}R_{t}\alpha_{t}\\ +(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}-S_{t}m_{t})^{\top}\bar{Q}_{t}(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}-S_{t}m_{t})\Big)\mathrm{d}t\\ +\frac{1}{2}\Big(\big(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{T}\big)^{\top}QX^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{T}\\ +\big(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{T}-Sm_{T}\big)^{\top}\bar{Q}\big(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{T}-Sm_{T}\big)\Big)\bigg].$ (2)
@@
It is customary in LQ MFG theory, that one does not try to solve the fixed point problem with respect to the flow of probability measures directly, but rather solves the adjoint fixed point problem with respect to the corresponding mean flows. Of course, in presence of the appropriate uniqueness properties, the two fixed point problems are equivalent. In our setting, the mean flow formulation is more beneficial since a) it leads to a natural extension of the classical theory to a rough setting and b) it does not make use of measure-valued paths/flows and avoids associated technical considerations.
Our results can be seen as extensions of [19, Thm 3.34], also Proposition 3.35 therein, on global solutions.
A novel feature or our analysis is to exploit the affine-linear rough noise structure of the agent dynamics (1) to rewrite it in a mild/Volterra form. Our first main result, stated here in a simplified form, establishes the equivalence between Volterra and rough MFGs; the precise formulation and proof follows from the combination of Theorems 4.5, 4.8 and 4.12 below.
@@
$\left\{\begin{aligned} &{X}^{m,\xi}_{t}=\mathcal{K}(t,0)\xi+M(m)_{t}\\ &\quad\quad\quad+\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}{X}^{m,\xi}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}Y^{m,\xi}_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}(t,s)\Sigma_{s}\mathrm{d}W_{s},\\ &Y^{m,\xi}_{t}=\mathcal{K}(T,t)^{\top}\big((Q+\bar{Q})X_{T}^{m,\xi}-\bar{Q}Sm_{T}\big)\\ &\quad\quad\quad+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{m,\xi}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{m,\xi}_{s}\Big)\mathrm{d}s-\int_{t}^{T}\mathcal{K}(s,t)^{\top}Z^{m,\xi}_{s}\mathrm{d}W_{s}.\end{aligned}\right$.
@@
A detailed version of this theorem is contained in Section 5 below as Theorem 5.3.
Finally, from the analysis conducted in Section 6, with respect to the problem discussed in the end of Section 1.1 we get [^4] See Theorem 6.7 for details.
Theorem 1.5.
Let $\xi\in\mathbb{R}^{d}$ and $_{t}+\bar{Q}_{t}-\bar{Q}_{t}S_{t}\succeq 0, Q+Q$$\bar{Q}-\bar{Q}S\succeq 0$. Under some specific technical assumptions [^5] See assumptions (S2), (S3) and (S6-R) below., there exist $(\overline{m^{*}},(\overline{m^{*}})^{\prime})\in\overline{\mathscr{M}}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$ and $\overline{\alpha^{*}}\in\overline{\mathcal{A}}$ such that
$\mathbb{E}\left[\int_{0}^{T}\ell^{\overline{m^{*}}}_{t}\big(\overline{X}^{\overline{\alpha^{*}}}_{t},\overline{\alpha^{*}}_{t}\big)\mathrm{d}t+\Phi^{\overline{m^{*}}}\big(\overline{X}^{\overline{\alpha^{*}}}_{T}\big)\bigg|\mathcal{F}^{\perp}_{T}\right]\\ =\left(\textrm{inf}_{\alpha\in\mathcal{A}}\mathbb{E}^{\prime}\!\left[\int_{0}^{T}\ell^{m(\boldsymbol{\eta})}_{t}\big({X}^{\boldsymbol{\eta}}_{t},\alpha_{t}\big)\mathrm{d}t+\Phi^{m(\boldsymbol{\eta})}\big({X}^{\boldsymbol{\eta}}_{T}\big)\right]\right)\Bigg|_{\boldsymbol{\eta}=\textbf{W}^{\perp}(\omega^{\prime\prime})}$
while
1.3 Main contributions and discussion
  1. 1.
    To the best of our knowledge, this work is, next to [9], the first to formulate MFG using rough paths [^6] The rough path framework allows for robustness (as can be seen by the results of Section 5) and great flexibility. as model for common noise, as was envisioned in [18]. As is standard in control theory, mean-field control and mean-field games, the linear-quadratic case serves as important model case and benchmark. Global well-posedness then hinges on non-explosion for new classes of rough Riccati equations. The work [9], which appeared shortly after our first version on the arXiv, extends work of Bo et al. [11] from the Poisson to the Brownian case, and employs relaxed solutions and compactness methods to obtain rough MFG existence (their results do not overlap with ours).
  2. 2.
    We introduce a Volterra-type (or mild) formulation, a priori independent of rough path considerations, that strikes a balance between a direct (rough stochastic) analysis of the mean-field game and a “fully flow transformed” equation, often considered as non-intrinsic. The control problem for the individual agent is then uniquely solvable via the stochastic maximum principle (SMP). The same approach should also work for non-linear $b,\Sigma$, under suitable strict convexity assumptions on the Hamiltonian; the advantage of (uncontrolled) affine rough drift is that is does not affect the convexity properties of the Hamiltonian, which is equally felt when switching to the mild formulation. From a technical perspective, the mild formulation streamlines certain arguments related to moment conditions in rough Itô calculus, as we explain in detail in Remark 3.8.
  3. 3.
    When $\Sigma$ is not invertible the standard conditions for existence of (global, strong) solutions of MFG (in absence of rough drift) via SMP are rather stringent, see e.g. [19, Assumption (SMP), p. 276] and [20, Assumptions (Sufficient SMP in Random Environment), p. 99, (MFG with a Common Noise SMP), p. 210], but very important. Hence, by smoothly approximating $\boldsymbol{\eta}$ and absorbing the rough integral in the drift term, one can see that assuming affine dependence on state in the rough integral is naturally aligned with the aforementioned sets of conditions. Furthermore, when rough common noise ceases to be affine in the agent’s state [^7] Thus, the mild formulation is no more appropriate., while one expects the SMP to still yield necessary optimality conditions for the representative agent’s control problem, she should not expect to give sufficient conditions for global solutions of the problem. This was demonstrated in a deterministic setting (absence of idiosyncratic noise) in [25] via a rough Pontryagin’s type maximum principle (PMP) and was noted by the authors in [25, Remarks 9, 25]. That is mainly because the (uncontrolled) sign of $\dot{\eta}$” destroys convexity properties of the Hamiltonian. Nevertheless, by employing rough FBSDE machinery one can instead aim for local well-posedness results, or by compactness methods for (global) existence, without uniqueness, as done in [9].
  4. 4.
    We establish, for the first time, well posedness of rough McKean-Vlasov FBSDEs, thanks to progress on rough Riccati equations, while demonstrating that the latter can characterize rough MFG solutions. Along the way, we also show that rough FBSDEs can be used to naturally extend the stochastic maximum principle to rough stochastic optimal control problems.
  5. 5.
    We exhibit the stochastic meaning of rough MFG, under independent Brownian randomization of the rough path. Indeed, this yields a novel rough MFG version of what several authors (in a non-mean field, purely control setting) studied on the name of pathwise stochastic control [50, 49, 12, 22, 29, 42].
@@ 1.4 Structure of the paper
After discussing frequently-used notation and some preliminaries on rough paths and rough (stochastic) integration at Section 2, we begin our approach to rough LQ MFG at Section 3. There, we develop the stochastic maximum principle in a Volterra setting by avoiding completely any mention of rough paths. Then, at Section 4 we connect the Volterra and rough settings by delicate use of Duhamel’s principle, rigorously proved with a rough product formula. Thus, we obtain the aforementioned existence and uniqueness results for the Volterra and Rough FBSDEs, while also reducing the problem of solutions for the Volterra and rough LQ MFG problems to their deterministic counterparts. Next, we solve the deterministic forward-backward rough differential equation (FBRDE) by following the classical approach and imposing more general assumptions to account for the presence of the rough path. In Section 5, we establish stability estimates and local-Lipschitz continuity of the Itô-Lions-Lyons map with respect to rough common noise and initial conditions. Finally, Section 6 is devoted to randomization of the common noise and connection to the (conditional) stochastic problem that was discussed at the end of Section 1.1. Auxiliary results on rough SDEs, accompanied by short proofs, are relegated to Appendix A.
@@ 2.2 Rough path theory
    The space of $\gamma$-Hölder continuous rough paths is denoted $\mathscr{C}^{\gamma}([0,T];\mathbb{R}^{p})$ and is equipped with the following rough path norm and rough path metric: for any $\boldsymbol{\eta}=(\eta,\eta^{(2)}),\bar{\boldsymbol{\eta}}=(\bar{\eta},{\bar{\eta}}^{(2)})\in\mathscr{C}^{\gamma}([0,T];\mathbb{R}^{p}),$
@@
We shall always endow $\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$with the rough path metric $\rho_{\gamma}$. Moreover, we may write $\rho_{\gamma}(\boldsymbol{\eta})$ instead of $\rho_{\gamma}(\boldsymbol{\eta},0)$.
@@ 3. $(\text{iii})$
    In this case we write $(Z,Z^{\prime})\in\textbf{D}_{\eta}^{\beta,\beta^{\prime}}\mathbb{L}^{k}([0,T];V)$ and $(Z,Z^{\prime})\in\mathbf{D}_{\eta}^{2\beta}\mathbb{L}^{k}([0,T];V)$ if $\beta=\beta^{\prime}$. We equip the space $\mathbf{D}_{\eta}^{\beta,\beta^{\prime}}L^{k}$ with the norms
    $\llbracket(Z,Z^{\prime})\rrbracket_{\eta,\beta,\beta^{\prime},k}$ $:=\left\lVert Z_{0}\right\rVert_{k}+\lVert\delta Z\rVert_{\beta,k}+\left\lVert Z^{\prime}_{0}\right\rVert_{k}+\lVert\delta Z^{\prime}\rVert_{\beta^{\prime},k}+\lVert\mathbb{E}_{\bullet}R^{Z}\rVert_{\beta+\beta^{\prime},k},$
    $\|(Z,Z^{\prime})\|_{\eta,\beta,\beta^{\prime},k}$ $:=\left\lVert Z_{0}\right\rVert_{k}+\lVert\delta Z\rVert_{\beta,k}+\left\lVert Z^{\prime}_{0}\right\rVert_{k}+\lVert\delta Z^{\prime}\rVert_{\beta^{\prime},k}+\lVert R^{Z}\rVert_{\beta+\beta^{\prime},k}.$
@@
    $\llbracket(Z,Z^{\prime});(\bar{Z},\bar{Z}^{\prime})\rrbracket_{\eta,\bar{\eta},\beta,\beta^{\prime},k}:=\|\delta(Z-\bar{Z})\|_{\beta,k}+\|\delta(Z^{\prime}-\bar{Z}^{\prime})\|_{\beta^{\prime},k}+\|Z_{0}-\bar{Z}_{0}\|_{k}$
@@
    $\|(Z,Z^{\prime});(\bar{Z},\bar{Z}^{\prime})\|_{\eta,\bar{\eta},\beta,\beta^{\prime},k}$ $\penalty 10000\ :=\|\delta(Z-\bar{Z})\|_{\beta,k}+\|\delta(Z^{\prime}-\bar{Z}^{\prime})\|_{\beta^{\prime},k}$
@@
We emphasize that in the case where $(Z,Z^{\prime})\in\mathbf{D}_{{\eta}}^{\beta,\beta^{\prime}}\mathbb{L}^{k}$ is deterministic then immediately we get $(Z,Z^{\prime})\in\mathscr{D}_{\eta}^{\beta,\beta^{\prime}}\mathbb{L}^{\infty}.$ The space of deterministic controlled rough paths is denoted by $\mathscr{D}_{\eta}^{\beta,\beta^{\prime}}$, or $\mathscr{D}_{\eta}^{\beta,\beta^{\prime}}([0,T];V)$ when we need to emphasize the codomain $V.$ Furthermore, in the deterministic setting we abbreviate the notation and we simply write $\|(\cdot,\cdot)\|_{\eta,\beta,\beta^{\prime}},\|(\cdot,\cdot);(\cdot,\cdot)\|_{\eta,\bar{\eta},\beta,\beta^{\prime}}$.
@@
$\mathop{\mathrm{ess\,sup}}_{t,\omega}|Z_{t}|\vee\|(Z,Z^{\prime})\|_{{\eta},\beta,\beta^{\prime},\infty}\leq K\quad\mbox{and}\quad\|z_{0}\|_{k}\vee\|z^{\prime}_{0}\|_{k}\leq K_{0}$
@@
The next result is from [13, Thm 3.11] in case $\beta^{\prime}=\beta$, and [42, Thm 3.5] otherwise. It establishes well-posedness of (1).
@@
Let $p\in\mathbb{N},$ $\gamma\in(\frac{1}{3},\frac{1}{2})$ and $0<\beta^{\prime}\leq\beta\leq\gamma$ such that $\gamma+\beta+\beta^{\prime}>1$, and $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$. Additionally, let $A,B,C:[0,T]\rightarrow\mathbb{R}^{d\times d}$ bounded and Borel measurable on $[0,T]$, $\big({A}^{1},(A^{1})^{\prime}\big),\big({C}^{1},(C^{1})^{\prime}\big)\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big),(m,m^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ deterministic controlled rough paths with respect to $\eta$}, $\xi\in\bigcap_{k=2}^{\infty}\mathbb{L}^{k}(\mathcal{F}_{0},\mathbb{R}^{d})$, and $\alpha(\omega,t),\Sigma(\omega,t)$ progressively measurable processes such that $\sup_{t\in[0,T]}\mathbb{E}[|\Sigma_{t}|^{k}]+\sup_{t\in[0,T]}\mathbb{E}[|\alpha_{t}|^{k}]<\infty$ for every $k\in[2,\infty)$. Then, there exists a unique solution $X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}$ of